# USM: Unbiased Survey Modeling for Limiting Negative User Experiences in Recommendation Systems

### v1

@@ Abstract.
Reducing negative user experiences is essential for the success of recommendation platforms. Exposing users to inappropriate content could not only adversely affect users’ psychological well-beings, but also potentially drive users away from the platform, sabotaging the platform’s long-term success. However, recommendation algorithms tend to weigh more heavily on positive feedback signals due to the scarcity of negative ones, which may result in the neglect of valuable negative user feedback. In this paper, we propose an approach aimed at limiting negative user experiences. Our method primarily relies on distributing in-feed surveys to the users, modeling the users’ feedback collected from the survey, and integrating the model predictions into the recommendation system. In addition, we strive to resolve two problems typically encountered in modeling user survey feedback: (a) We address the Response Bias by applying a survey-submit model; (b) We address the Exposure Bias by adopting a pseudo label distillation training framework. The A/B testing results indicate a reduction in survey sexual rate and survey inappropriate rate, ranging from -1.44% to -3.9%. Additionally, we compared our methods against an online baseline that does not incorporate our approach. The results indicate that our approach significantly reduces the report rate and dislike rate by 1% to 2.27% compared to the baseline, confirming the effectiveness of our methods in enhancing user experience. After we launched the survey model based our approach on our platform, the model is able to bring reductions of 1.75%, 2.57%, 2.06% on reports, dislikes, survey inappropriate rate, respectively.
@@ 1. Introduction
In this paper, we develop an in-feed survey model that utilizes user and item features as inputs, with survey responses serving as labels. We subsequently integrate this model into the recommendation system to optimize the users’ FYFs. This approach aims to enhance user satisfaction with the videos recommended by the platform, ultimately leading to improvement of daily active users (DAU) and corresponding TnS metrics. However, due to the sparse distribution of in-feed survey and users’ different willingness to submit survey, we face significant challenges in terms of exposure bias and response bias during model training. Correspondingly, we propose two approaches to address these challenges.
The main contribution of this work can be briefly summarized as: 1)1)1 ) We distributed two types of surveys within FYF and utilized the survey responses to build the in-feed survey model, achieving significant gains in daily active users (DAU) and corresponding content issue metrics. 2)2)2 ) We developed a survey-submit model and used the estimated submission rates to weight the samples in the in-feed survey model, thereby resolving the issue of response bias. 3)3)3 ) We developed a pseudo label distillation framework to use a teacher model to generate pseudo survey labels for all video view samples, and subsequently utilize these labels to train a student model for online serving, thereby effectively mitigating the issue of exposure bias.
@@ 2. Related Work
  • Exposure Bias. Exposure Bias occurs when users are only exposed to a subset of specific items, meaning that unobserved interactions do not necessarily indicate a negative preference. Previous studies have explored various aspects of data exposure: 1) Exposure is influenced by the policies of earlier recommendation systems, which dictate which items are shown to users (Liu et al., 2020). 2) Users actively search for items of interest, and their selection process significantly affects exposure. This means that highly relevant items are more likely to be presented (Wang et al., 2016), leading to what is often referred to as “user-selection bias.” 3) Popular items tend to receive more visibility, which introduces a form of ”popularity bias,” another variant of “exposure bias” (Zheng et al., 2021). In order to facilitate readers and prevent concept confusion, we use the unified standard definition, “exposure bias”, throughout this paper, rather than the separated definitions of the aforementioned factors. In personalized survey modeling, surveys are randomly distributed to users at a low frequency. As a result, the results of users’ survey responses for many items are not observable. This paper examines this exposure bias from the perspective of survey modeling and proposes a solution based on pseudo label distillation.
@@ 3. Approaches
In this section, we will first outline the structure of the baseline in-feed survey model and its application in recommendation systems. To tackle the challenges of response bias and exposure bias in survey modeling, we will introduce two solutions: the development of a survey-submit model and the implementation of a pseudo label distillation framework.
@@ 3.1. In-feed Survey Model
The structure of the in-feed survey model is depicted in the Fig 3. The architecture of the model is structured as a multi-head framework, in which each head estimates the probability of a user selecting a specific option. The backbone utilizes universal embeddings for users, items, and authors, along with additional raw attribute features of users and items as inputs, and consists of three fully connected (FC) layers with output dimensions of {512, 256, 128}. Subsequently, we design a multi-head architecture, where each head comprises three FC layers with output dimensions of {64, 16, 1}. In addition, we enhance the backbone of the model by incorporating the Learning Hidden Unit Contributions (LHUC) Module and the Squeeze-and-Excitation (SE) Module to facilitate feature cross-interaction and self-attention extraction, both of which have been substantiated as effective in numerous studies (Swietojanski et al., 2016; Hu et al., 2018; Wu et al., 2019). The LHUC module identifies and selects critical universal features that are not integrated into the backbone network, specifically user_id, item_id, and author_id. These features are subsequently applied to the inputs of the three fully connected layers, thereby facilitating feature interaction. This methodology enhances the model’s capacity to learn implicit information derived from both the inputs of the fully connected layers and the original attributes of users and items. The SE Module is applied to the output of the final FC layer of the backbone, employing self-attention mechanisms to dynamically weight and recalibrate channel-wise features adaptively.
@@ 3.2.2. Survey-Submit Model
3.3. Exposure Bias & Pseudo Label Distillation
3.3.1. Problem Formulation
Exposure bias is a common issue in recommendation models. This bias occurs because the training data for the model is based only on data from items or users that have been exposed, which leads to an under-representation of items or users that are not as well-exposed. In the context of survey modeling, the training data includes only user-item pairs for which surveys have been distributed.
To minimize disruption for users, we controlled the frequency of survey distributions, which led to significant exposure bias in the model. On one hand, the model has a limited number of labeled positive and negative examples for user-item pairs, accounting for less than one ten-thousandth of the total video exposures. This lack of training data restricts the model from fitting only local distributions. On the other hand, each user receives only one survey sample within a short time frame, further limiting the granularity of personalization in the modeling process.
In our baseline model, we focused on generalized features to enhance the model’s ability to generalize beyond the training data, alleviating exposure bias and achieving coarse-grained personalization.Moreover, we discovered that by fully utilizing exposure data, we could reduce exposure bias while significantly increasing our training sample size. To address this, we propose a pseudo label distillation approach that assigns pseudo labels to all exposed videos for which surveys have not been distributed.
3.3.2. Overall Structure
The pseudo label distillation approach, as illustrated in Fig 4, consists of two sample streams. The first stream comprises samples from survey exposures, which collect user-submitted responses and thus represent labeled data. The second stream includes all video exposure data.
It is clear that the number of samples in the first stream is significantly smaller than that in the second stream. We train a teacher model using the first sample stream, which then makes predictions on the second sample stream to generate pseudo labels. The video exposure data, processed through a pseudo label filter, is combined with the first sample stream to guide the training of the student model.
Figure 4. Pseudo Label Distillation Structure. The teacher model is trained on labeled samples and generates pseudo labels for unlabeled samples to guide the training of the student model. The features of the teacher and student models are largely consistent, with the only distinction being that the teacher model possesses posterior behavioral
3.3.3. Teacher Model
The structure of the teacher model is consistent with the baseline model. Input features consist of item features, user features, contextual features, and posterior behavior features. The posterior behavior features refer to the user’s feedback on the video after it is exposed, such as whether or not to like or share the video. These feedbacks provide additional information that can enhance the teacher model’s accuracy. To better align with inappropriate survey tasks, we added user behaviors that express negative experiences as posterior behavior features, such as dislikes and reports. It is important to note that posterior behavior features cannot be obtained prior to the exposure of videos; therefore, they are not available during the online service phase and can only be utilized within the teacher model.
The format of the model loss function is presented as follows, where $y_{u,i}$ represents the responses provided by user $u$ after receiving the survey for item $i$, and $p_{u,i}$ denotes the estimated score produced by the teacher model.
(7) $loss=CE(p_{u,i},y_{u,i})$
3.3.4. Student model
The training data consists of two parts: The first part consists of samples without survey exposure, where the teacher model estimates a continuous value $\tilde{y}_{u,i}$ as a pseudo label. Given that the teacher model is not entirely accurate, it may generate numerous erroneous pseudo labels, leading to noisy training. Following the method described in reference (Rizve et al., 2021), we established a threshold for both positive and negative examples within the pseudo labels to exclude certain samples whose predicted results fall within an uncertain intermediate range, where the teacher model exhibits lower confidence. Additionally, we continue to utilize soft labels to enable the student model to learn the subtle differences between different predicted values. Let $g_{u,j}$ indicate whether the pseudo label for the sample is used for training the model:
(8) $g_{u,i}=\begin{cases}1&\text{if }\tilde{y}_{u,i}<\tau_{n}\text{ or }\tilde{y}_{u,i}>\tau_{p}\\ 0&\text{if }others\end{cases}$
where $\tau_{p}$ and $\tau_{n}$ represent the thresholds for selecting pseudo labels for positive and negative examples, respectively. The second part of the training data is samples with survey exposure, labeled as $y_{u,i}$.
The structure of the student model can be simpler than that of the teacher model. Since the teacher model in this study is not complex, it does not significantly increase online latency; therefore, the student model uses the same network architecture. However, the student model cannot include posterior behavior features, as it is required to operate in an online service phase. Incorporating posterior behavior features would lead to inconsistencies between offline and online predictions.
The format of the model loss function is given by Eq. 9, where $\tilde{y}_{u,i}$ is the estimated score from the teacher model on unlabeled samples, and $\alpha$ is used to adjust the weight of labeled samples.
(9) $loss=CE(p_{u,i},\tilde{y}_{u,i})+\alpha*CE(p_{u,i},y_{u,i})$
@@ 4.1.1. AUC and Calibration
For labeled samples, we utilize AUC (Area Under the Curve) to assess the offline model’s ability to rank positive examples ahead of negative examples. In the case of teacher-student distillation, where the supervisory signal from the teacher model is represented by continuous predicted values, we employ regAUC to evaluate the student model’s ranking capability. Additionally, we use calibration values to measure the discrepancy between the model’s average predictions and the actual event occurrence rates. The calculation method for calibration values is as follows:
@@ 4.1.2. Neg/Pos-Feedback UAUC
To address the issue of exposure bias, we introduced a pseudo label distillation approach to model users’ negative experiences more accurately when surveys are not presented. Because there are no explicit labels for unexposed surveys, it is impossible to directly calculate the AUC to assess the modeling accuracy of these sample distributions.
As mentioned earlier, there is a significant correlation between user reports (or dislikes) and the responses indicating ”inappropriate” in the surveys, and reports (or dislikes) buttons are exposed to all users. Therefore, we use reports and dislikes as implicit ground truth to evaluate the accuracy of the student model. We first calculate the Nfb (Negative-Feedback) AUC for each user, as
@@ 4.2. Survey-Submit Model
4.3. Pseudo Label Distillation
To ensure that each step of the process meets expectations, we evaluate both the teacher model and the student model separately.
As the teacher model is not deployed for online serving, we conducted only offline AUC evaluations. The introduction of posterior behavior features significantly enhanced the prediction accuracy of the teacher model. Specifically, we observed improvements of 1.8% in the satisfaction survey AUC, 1.6% in the sexual survey AUC, and 1.8% in the inappropriate survey AUC.
The student model is deployed for online serving and plays a crucial role in our system. Table 3 presents the online negative/positive-feedback UAUC for the student model. For all three predicted targets, the treatment group exhibits higher performance than the control group, indicating that pseudo label distillation more accurately models negative user feedback on videos where surveys have not been exposed.
Head Model UAUC
Satisfaction Survey Baseline 0.7615
Distillation 0.7749
Distillation + Response Debias 0.7771
Sexual Survey Baseline 0.6363
Distillation 0.6442
Distillation + Response Debias 0.6491
Inappropriate Survey Baseline 0.6581
Distillation 0.6813
Distillation + Response Debias 0.6861
Table 3. Neg/Pos-Feedback UAUC. Satisfaction Survey is evaluated by Pos-Feedback UAUC, Sexual Survey and Inappropriate Survey is evaluated by Neg-Feedback UAUC
As illustrated in Fig 6, we observed significant improvements in negative feedback metrics during our online A/B testing. Specifically, in survey responses, we noted a 0.47% decrease in the survey sexual rate and a 1.62% reduction in the survey inappropriate rate. Regarding user-initiated negative feedback behaviors, the dislike rate decreased by 2.27%, while the report rate declined by 1.47%. Through long-term reversal experiments, we found that these improvements led to a 0.01% increase in long-term user retention, further validating the effectiveness and lasting impact of our approach.
To further validate the effectiveness of our approaches, we combined pseudo label distillation with the survey-submit model. Specifically, we used the predicted scores generated by the survey-submit model to reduce response bias in the pseudo labels of the student model, addressing both response bias and exposure bias. After deploying the trained model on the online platform, we observed that its online UAUC was higher than that of the model that relied solely on the pseudo label distillation method, as indicated by online predictive accuracy metrics. For more details, please refer to Table 3, which illustrates the effectiveness of our approach. We also witnessed notable enhancements in negative feedback metrics during our online A/B testing when compared with the model employing pseudo label distillation alone. As illustrated in Fig 6, the survey sexual rate decreased by 2%, and the survey inappropriate rate was reduced by 1.95%.
After successfully achieving positive online UAUC and favorable results from our online A/B testing, we launched the survey model based on our approaches on our platform. In the months following the launch, we saw a 2.06% decrease in survey inappropriate rate, which was determined through the survey results submitted by users on our platform. Additionally, we monitored other negative action metrics and observed positive trends in these areas as well. Specifically, the report rate saw a decrease of 1.75%, suggesting that users are encountering fewer issues warranting complaints. Furthermore, the dislike rate also improved, with a decrease of 2.57%, indicating greater overall satisfaction with the content presented on our platform.These outcomes not only reflect the effectiveness of the corrective measures we have implemented but also highlight our commitment to continuous improvement. We remain dedicated to monitoring these metrics and further refining our strategies to enhance user engagement and satisfaction.
Figure 6. Negtive Feedback Metric Reduction using Pseudo Label Distillation and Response Debias

### v3

@@ Abstract.
Reducing negative user experiences is essential for the success of recommendation platforms. Exposing users to inappropriate content could not only adversely affect users’ psychological well-beings, but also potentially drive users away from the platform, sabotaging the platform’s long-term success. However, recommendation algorithms tend to weigh more heavily on positive feedback signals due to the scarcity of negative ones, which may result in the neglect of valuable negative user feedback. In this paper, we propose an approach aimed at limiting negative user experiences. Our method primarily relies on distributing in-feed surveys to the users, modeling the users’ feedback collected from the survey, and integrating the model predictions into the recommendation system. We further enhance the baseline survey model by integrating the Learning Hidden Unit Contributions module and the Squeeze-and-Excitation module. In addition, we strive to resolve the problem of response Bias by applying a survey-submit model; The A/B testing results indicate a reduction in survey sexual rate and survey inappropriate rate, ranging from -1.44% to -3.9%. Additionally, we compared our methods against an online baseline that does not incorporate our approach. The results indicate that our approach significantly reduces the report rate and dislike rate by 1% to 2.27% compared to the baseline, confirming the effectiveness of our methods in enhancing user experience. After we launched the survey model based our approach on our platform, the model is able to bring reductions of 1.75%, 2.57%, 2.06% on reports, dislikes, survey inappropriate rate, respectively.
@@ 1. Introduction
In this paper, we develop an in-feed survey model that utilizes user and item features as inputs, with survey responses serving as labels. We further enhance the baseline survey model by integrating the Learning Hidden Unit Contributions (LHUC) module and the Squeeze-and-Excitation (SE) module, which are utilized to selectively incorporate both universal and critical features, and to capture self-attention information, respectively. We subsequently integrate this model into the recommendation system to optimize the users’ FYFs. This approach aims to enhance user satisfaction with the videos recommended by the platform, ultimately leading to improvement of daily active users (DAU) and corresponding TnS metrics. However, due to the sparse distribution of in-feed survey and users’ different willingness to submit survey, we face significant challenges in terms of exposure bias and response bias during model training. Correspondingly, we propose two approaches to address these challenges.
The main contribution of this work can be briefly summarized as: 1)1)1 ) We distributed two types of surveys within FYF and utilized the survey responses to build the in-feed survey model, achieving significant gains in daily active users (DAU) and corresponding content issue metrics. 2)2)2 ) We developed a survey-submit model and used the estimated submission rates to weight the samples in the in-feed survey model, thereby resolving the issue of response bias.
@@ 3. Approaches
In this section, we will first outline the structure of the baseline in-feed survey model and its application in recommendation systems. To tackle the challenges of exposure bias in survey modeling, we will introduce a survey-submit model.
@@ 3.1. In-feed Survey Model
The structure of the in-feed survey model is depicted in the Fig 3. The architecture of the model is structured as a multi-head framework, in which each head estimates the probability of a user selecting a specific option. The backbone utilizes universal embeddings for users, items, and authors, along with additional raw attribute features of users and items as inputs, and consists of three fully connected (FC) layers with output dimensions of {512, 256, 128}. Subsequently, we design a multi-head architecture, where each head comprises three FC layers with output dimensions of {64, 16, 1}.
@@
3.2. LHUC Module & SE Module
We enhance the backbone of the model by incorporating the Learning Hidden Unit Contributions (LHUC) module and the Squeeze-and-Excitation (SE) module to facilitate feature cross-interaction and self-attention extraction, both of which have been substantiated as effective in numerous studies (Wu et al., 2019a, 2021; Zhu et al., 2020; Wu et al., 2019b). The structures of the LHUC and SE modules are shown in Fig 4.
3.2.1. LHUC Module
The LHUC module (Swietojanski et al., 2016) aims at enhancing model performance by selectively incorporating both universal and critical features, improving feature interaction, and enabling the model to learn implicit information from both the original attributes of users and items and the inputs to the fully connected layers, ultimately improving personalized predictions and overall accuracy.
Universal features primarily encompass fundamental information about users, videos, and authors, such as user_id, item_id, and author_id. Critical features, on the other hand, are determined through the calculation of feature importance, which quantifies the relevance of each feature in the context of the model’s task. Feature importance is computed using a masking technique, where, during model evaluation, the corresponding feature is set to zero. The importance of a feature is inversely proportional to the decrease in model AUC: a larger reduction in AUC after masking a feature indicates a higher level of importance for that feature, as it signifies a more significant contribution to the model’s predictive performance. The top-3 most important features selected for the LHUC module inputs are language, region, and device, which aligns with our intuition. All three features are closely related to the user’s context and significantly influence user preferences and behaviors. Language reflects the user’s communication preferences, region indicates geographic location and potential cultural influences, and device offers insights into the user’s platform of interaction. These contextual factors play a significant role in shaping user behavior, making them crucial for personalized predictions and recommendations.
The structure of the LHUC module is shown in Fig 4(a). The module employs input embeddings to generate three output embeddings through a series of fully connected layers, with dimensions of {512, 256, 128}, which are precisely aligned with the output dimensions of the multi-FC layers in the backbone network. The LHUC output embeddings $LHUC_{i}$ are channel-wise multiplied with the outputs of the multi-FC layers $FC\_\text{origin}_{i}$ to produce the final outputs $FC\_\text{final}_{i}$, enabling effective feature modulation, as shown in Eq.4:
(4) $FC\_\text{final}_{i}=LHUC_{i}\cdot FC\_\text{origin}_{i}$
, where $i$ corresponds to the number of layers in the multi-FC network.
3.2.2. SE Module
The SE module (Hu et al., 2018) is applied to the output of the final FC layer of the backbone, employing self-attention mechanisms to dynamically weight and recalibrate channel-wise features adaptively.
The structure of the SE module is shown in Fig 4(b). The SE module initially applies a global pooling operation to squeeze the input along the channel dimension, producing a statistic $\mathbf{s}\in\mathbb{R}^{C}$, where $C$ represents the channel dimension. Subsequently, the SE module utilizes two fully connected layers to fully capture channel-wise dependencies $e$, as expressed by Eq.5:
(5) $e=\sigma(W_{2}\delta(W_{1}\mathbf{s}))$
, where $\delta$ refers to the ReLU active function (Nair and Hinton, 2010), $\sigma$ refers to the Sigmoid active function (Rumelhart et al., 1986), $W_{1}\in\mathbb{R}^{C/n\times C}$ and $\quad W_{2}\in\mathbb{R}^{C\times C/n}$. We set $n=4$ to optimize the trade-off between computational efficiency and representation capacity, ensuring reduced resource consumption of the SE module while maintaining the integrity of inter-channel dependencies. The final output of the SE module is produced by rescaling the input using the activation $e$:
(6) $output=F_{\text{scale}}(input,e)=e\cdot input$
, where $F_{\text{scale}}$ refers to channel-wise multiplication.
(a) LHUC module
(b) SE module
Figure 4. LHUC & SE module detail structure. (4(a)) The LHUC module utilizes the original input features to generate three output embeddings, which are multiplied with the outputs of the backbone FC layers;. (4(b)) The SE module performs global pooling for squeezing, followed by two FC layers to extract self-attention, and multiplies it with the input for excitation.
@@ 4.1.1. AUC and Calibration
For labeled samples, we utilize AUC (Area Under the Curve) to assess the offline model’s ability to rank positive examples ahead of negative examples. Additionally, we use calibration values to measure the discrepancy between the model’s average predictions and the actual event occurrence rates. The calculation method for calibration values is as follows:
@@ 4.1.2. Neg/Pos-Feedback UAUC
When the predicted models scores are used in the ranking phase, we place greater emphasis on the user-weighted average area under the curve (UAUC) metric. The UAUC is defined as an average of the area under the curve (AUC) for each user, aimed at assessing the model’s ranking ability for different candidate items specific to individual users. However, in survey scenarios, the sparsity of feedback makes it challenging to collect responses from the same user with respect to multiple items, resulting in an inability to compute the UAUC metric for the survey behavior. However, we find a significant correlation between user reports (or dislikes) and responses indicating ”inappropriate” content in surveys. Furthermore, since report (or dislike) buttons are exposed to all users, this leads to a denser collection of implicit feedback, allowing us to gather preferences for reports and dislikes from the same user across different candidate items. Therefore, we utilize reports and dislikes as implicit ground truths to evaluate the accuracy of the student model. We first calculate the Nfb (Negative-Feedback) AUC for each user, as
@@ 4.1.4. Online Debiased A/B Testing
4.2. LHUC/SE Modules
The experimental results show that integrating the LHUC/SE modules improves model performance across various surveys, as illustrated in Table 3. The AUC improvements indicate more accurate predictions of survey responses. Additionally, the negative/positive feedback UAUC metrics reveal that the model can better distinguish between individual users’ positive and negative feedback on candidate items. This integration significantly enhances the model’s ability to process both types of feedback, highlighting its superiority over the baseline model.
We also observed significant improvements in negative feedback metrics during our online A/B testing. Specifically, in the survey responses, we noted a 0.70% decrease in the survey sexual rate and a 0.58% reduction in the survey inappropriate rate. Regarding user-initiated negative feedback behaviors, the dislike rate decreased by 0.81%. Through long-term reversal experiments, we found that these improvements led to a 0.01% increase in long-term user retention, further validating the effectiveness and lasting impact of our approach.
@@ 4.3. Survey-Submit Model
To further validate the effectiveness of our approaches, we combined LHUC and SE block with the survey-submit model. Specifically, we used the predicted scores generated by the survey-submit model to reduce response bias in survey model, addressing response bias. After deploying the trained model on the online platform, we observed that its online UAUC was higher than that of the model that relied solely on LHUC and SE block, as indicated by online predictive accuracy metrics. For more details, please refer to Table 3, which illustrates the effectiveness of our approach.
@@
Head Model AUC UAUC
Satisfaction Survey Baseline 0.7416 0.7000
LHUC 0.7720 0.7123
LHUC+Debias 0.7722 0.7173
Sexual Survey Baseline 0.7531 0.6678
LHUC 0.7956 0.6759
LHUC+Debias 0.7966 0.6810
Inappropriate Survey Baseline 0.7846 0.6990
LHUC 0.8040 0.7003
LHUC+Debias 0.8061 0.7051
Table 3. AUC and Neg/Pos-Feedback UAUC. Satisfaction Survey is evaluated by Pos-Feedback UAUC, Sexual Survey and Inappropriate Survey is evaluated by Neg-Feedback UAUC